OpenStudy (anonymous):

I have to turn this in before 10 tonight can someone help? Evaluate x/y for x = 2/3 and y = 5/6.

OpenStudy (anonymous):

I don't have a clue what this is.

OpenStudy (danjs):

Dividing by a fraction is the same as multiplying by the fraction flipped over...

OpenStudy (danjs):

\[\frac{ \frac{ 2 }{ 3 } }{ \frac{ 5 }{ 6 } } = \frac{ 2}{ 3 } * \frac{ 6 }{ 5 }\]

OpenStudy (anonymous):

I still don't understand. :(

OpenStudy (anonymous):

Ooooohh I got it thanks DanJS

OpenStudy (danjs):

ok, no prob, you just multiply the numerators and denominators together, and reduce if you can

OpenStudy (danjs):

12 / 15 = 4 / 5

OpenStudy (anonymous):

Yeah that's why I was having problems I forgot to reduce.

OpenStudy (danjs):

ahh, cool